OpenStudy (ashleyisakitty):

what are the foci of the ellipse given by the equation 25x^2+16y=400

random231 (random231):

can u write the general form of the ellipse from this equation?

OpenStudy (ashleyisakitty):

(x-h)^2/a^2+(y+k)^2/b^2?

random231 (random231):

yup!

random231 (random231):

dont forget the = 1!!!!!!

OpenStudy (ashleyisakitty):

I dont know what to plug in though :( h, a, k b?

random231 (random231):

do one thing:: divide both the sides of the equation by 400. and tell me the equation that you get!

OpenStudy (ashleyisakitty):

well 400/25=16 and 400/16 .. is 25. What do I do with these numbers?

random231 (random231):

so 16=a^2 and b^2=25

OpenStudy (ashleyisakitty):

(25-?)/16+(16+?)/25=1?

random231 (random231):

no no!!!